OpenStudy (anonymous):

1. x – 3 > –9 2. –2x + 1 ≤ –11 3. 10 < –3x + 1 4. 2(x + 5) > 8x – 8 5. –2(x – 3) ≥ 5 – (x + 3) 6. Explain, in complete sentences, when you would use an open circle or a closed circle, and when you would shade to the right or left, to graph an inequality on the number line. 7. Describe a real life scenario where inequalities are used and state the inequality. what i mainly need help on is how to graph them pls help?

OpenStudy (anonymous):

lets do the first one

OpenStudy (anonymous):

\(x-3 > -9\) start by adding 3 to both sides to get \(x\) on one side of the inequality to get \[x>-6\]

OpenStudy (anonymous):

well im pretty sure the answer to first one is X > -6..right? but i dont know how to graph that, which way does the line point to? i know its an open circle...

OpenStudy (anonymous):

now we read it to make sense out of it "\(x\) is GREATER than \(-6\)"

OpenStudy (anonymous):

so what kind of numbers are greater than \(-6\)? well, \(-5,-4,-3,..., 0,0,1,2,...\)

OpenStudy (anonymous):

ohhh! okay so it would be to the right. oh my god thank u !!!!!

OpenStudy (anonymous):

that is, number to the RIGHT of \(-6\) on the number line

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

make sure you know how to read your answer in english that will save you from making mistakes and always make \(x\) the subject, so even if you see an answer that looks like \(5<x\) you know to read it not as "5 is less than \(x\)" even though that is correct read it as "\( x\) is greater than 5"

OpenStudy (anonymous):

then in your head pick some numbers that work, that will let you know what direction to go in

OpenStudy (anonymous):

ok thank u i understand it so much better now!!

OpenStudy (anonymous):

want to try the next one? it is slightly more difficult

OpenStudy (anonymous):

well i have the answer to this one, but not the rest of the questions yet im pretty sure it's x ≥ 6 meaning it would then be a closed circle to the right also..? am i correct?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

yes that is correct, you should get \(x\geq 6\) which you represent as a closed circle because of the \(\leq\) and to the right. guess you got this good work

OpenStudy (anonymous):

i means closed circle because of the \(\geq\)

OpenStudy (anonymous):

yes thank u so much!!

OpenStudy (anonymous):

yw